OpenStudy (anonymous):

excuse me, but when you are solving an equation like like 5-x=4+x, *NOT LOOKING FOR ANSWER* when you subtract 4 from the right side does THAT number become negative? and if so, the left number is positive. if the number was 6 in the right, would it be -1? and what if the number on the left was also being subtracted (hence the equation becoming untrue) would that number also be a negative when subtracted by 4 ? THANKS

OpenStudy (anonymous):

when you subtract 4 from the right, there is no longer a 4 on the right, only an \(x\) when you subtract 4 from the left, you are subtracting it from the number on the left, namely 5 therefore you are doing the following arithmetic \(5-4\)

OpenStudy (unklerhaukus):

\[a=b+c\]\[a-c=b+c-c\]\[a-c=b\]

OpenStudy (anonymous):

if the equation was 5-6 however and when you are subtracting (not part of my question, confusing, next time ill slow down...) will that give us a negative or was the number that you subtracted become positive?

OpenStudy (anonymous):

the numbers do not "become" anything. you are doing whatever arithmetic needs to be done add, subtract, multiply, divide etc. whether the outcome is positive or negative depends entirely on the arithmetic being done

OpenStudy (anonymous):

i think what am trying to say is that our may be thinking about this in the wrong way your job is to isolate \(x\) and you do it doing arithmetic as needed

OpenStudy (anonymous):

oh, I see, only if the number Is negative does it become that, it was really confusing, because in an equation, like 5-x=x2-10 i thought when you subtracted the numbers became negative, hence you subtract a negative and a negative, THANK YOU!

OpenStudy (anonymous):

is this the equation \(5-x=x^2-10\) ?

OpenStudy (anonymous):

um, heh, well, it was actually something completly different, I think I can solve it myself now, really, again thanks a whole bunch SA! :D

OpenStudy (anonymous):

yw